OpenStudy (anonymous):

-3sqrt{12} + 3sqrt{3} + 3 sqrt{20} does it simplify to 9 sqrt {3} + 12 sqrt {5}

OpenStudy (anonymous):

no answer is\[-6\sqrt{2}+3\sqrt{3}+6\sqrt{5}\]

OpenStudy (anonymous):

How did you get\[-6\sqrt{2}\]

OpenStudy (anonymous):

sry it is actually \[-6\sqrt{3}\]

OpenStudy (anonymous):

Thanks this will help for tomorrows quiz